# Zero shot VLMs for hate meme detection: Are we there yet?

### v2

@@ 7 Results
vn vn 54.14 38.59 49.27 46.86 31.91 49.18 44.25 40.16 55.24 43.98 33.42 48.64
def vn 51.12 34.55 50.41 49.74 34.44 50.29 49.12 48.65 50.84 48.63 41.05 48.1
ocr vn 50.1 33.73 50.2 48.37 33.94 48.85 65.44 59.86 59.65 48.48 46.96 47.85
def + ocr vn 50.21 34.87 50.01 51.49 38.19 51.49 64.13 52.84 54.33 51.49 44.12 52.87
vn ex 48.38 38.06 49.05 50.35 35.2 50.4 46.63 41.27 41.27 44.84 44.21 44.69
def ex 49.68 33.19 50 51.43 49.85 51.56 46.88 46.77 51.04 50.35 50.27 50.41
ocr ex 49.12 34.41 49.75 47.39 47.34 47.56 65.42 55.19 55.91 49.5 45.17 48.7
def + ocr ex 53.06 44.37 52.72 54.39 52.52 53.97 65.6 51.25 53.61 54.09 49.82 53.4
InstructBLIP Flan-T5 xl
vn vn 50.2 33.78 50.2 56.67 48.6 56.79 64.12 41.88 50.09 51.27 36.32 50.07
def vn 50.2 35.13 50.2 59.9 56.67 59.9 64.97 39.38 50 51.27 33.89 50
ocr vn 51.2 39.22 51.2 55.9 47.75 55.9 65.16 40.21 50.4 51.27 33.89 50
def + ocr vn 52.6 42.33 52.6 52.1 39.59 52.1 65.54 42.52 51.18 51.98 36.68 50.63
vn ex 60.78 60.76 60.76 55.34 50.7 56.32 60.19 41.53 47.07 49.12 46.76 53.79
def ex 51 40.86 51 53.14 34.92 49.91 64.2 39.1 50 51.46 33.98 50
ocr ex 60 58.17 59.72 61.13 59.39 59.87 61.7 38.16 49.43 49.33 33.03 49.33
def + ocr ex 57.94 55.58 58.02 55.01 45.67 54.63 65.04 39.41 50 51.56 34.02 50
Table 2: Overall results - Accuracy, Macro-F1 and AUC-ROC score for 4 datasets and 5 models across 8 prompt variants per (model, dataset) combination. Greyed out cells signify that (model, prompt) combination is unable to classify for at-least 90% cases for the corresponding dataset. Best (model, prompt) combination per (model, dataset) combination is highlighted in light blue. Best (model, prompt) combination over each dataset is marked in yellow. in: Prompt input, out: Prompt output, vn: vanilla, def: definition, ocr: OCR text, ex: explanation.
@@
in out IDEFICS 9B LLaVA-1.5 13B LLaVA-1.5 7B I-BLIP V I-BLIP F
vn vn 43.46 55.47 35.26 NA 42.2
def vn NA 57.86 45.78 NA 45.36
ocr vn 52.34 53.7 41.02 40.14 42.38
def + ocr vn 40.4 missing62.03 56.84 NA 40.21
vn ex 40.52 55.89 41.95 NA NA
def ex 37.53 50.03 48.29 NA NA
ocr ex 49.84 59.26 39.83 NA NA
def + ocr ex 39.72 49.57 60.46 NA NA
Table 3: Leaderboard - Weighted macro F1 score for each (model, prompt) combination averaged across all 4 datasets. Overall best score is underlined and highlighted. Best scores across each prompt strategy are underlined. I-BLIP V: InstructBLIP Vicuna 7B, I-BLIP F: InstructBLIP Flan-T5-xl.
@@
In this section, we present the results of our experiments. In Table 2 we show the results for the four datasets across the five models. Each block in the table corresponds to a particular (model, dataset) combination and covers the results for eight prompt pattern combinations. Since we use the generation capability of VLMs for prediction, we observe that in some prompt variants, certain (model, prompt) combinations did not classify the input meme amongst the list_of_labels and diplomatically bypassed the query with an irrelevant answer. This led to a decrease in support to infer the results accurately grounded on correct labels in the dataset. In the table, we have greyed out the cases which did not generate a correct label for at least 90% of the data points. Examples of ambiguous outputs are provided in Appendix D.
@@
Overall results: From Table 2, we observe that InstructBLIP models are not able to correctly predict the labels out of list_of_labels and generate ambiguous answers for quite a large number of prompt variants. Their generated output did not conform with the expected output format specified in input prompt. We also observe that IDEFICS performs best with only OCR as input. Overall, LLaVA-1.5 13B emerges to be the best model with OCR and definition as input and vanilla output. It was also the most stable in terms of the responses generated out of all the VLMs considered across different prompt variants. LLaVA-1.5 7B, worked best with explanation as output, when the input prompt was definition and OCR text.
@@
Leaderboard : Since engineering solutions are always in the ‘quest for the best’, we propose a quantitative metric to organize the (model, prompt) combinations into a leaderboard. The idea is that the top combinations on this leaderboard should generalize well across the four datasets combined. For each prompt variant considered over all models, we calculate a weighted average macro F1 score depending on the number of samples in each of the datasets by the formulation: $\frac{\sum\limits_{\mathcal{D}}(f_{\mathcal{D}})*|\mathcal{D}|}{\sum\limits_{\mathcal{D}}|\mathcal{D}|}$. Here $f_{\mathcal{D}}$ is the macro F1 for the dataset $\mathcal{D}$. The results are shown Table 3. We did not consider those prompt variants in our calculation which did not produce results for at least $90\%$ of the data points i.e., the grey entries in Table 2. For such cases, we mark NA in Table 3. Based on the above results, we conclude LLaVA-1.5 13B to be the best model with definition and OCR text as input and vanilla as output. Further, in 7 out of 8 prompt variants LLaVA-1.5 13B outperforms all other models across the datasets combined. The only variant where LLaVA-1.5 7B beats LLaVA-1.5 13B is definition and OCR text as input and explanation as output.
8 Error analysis
@@
In the previous section we found that LLaVA-1.5 13B (with definition + OCR text as input and vanilla as output) is at the top of the leaderboard. We therefore investigate the cases of misclassification for this setting by comprehensively evaluating a total of 799 misclassified memes across considered datasets; 202 from FHM, 321 from MAMI and 276 from HARM P+C datasets. In particular we attempt to obtain an explanation of parts in the meme that confuses the model resulting in the mispredictions (section 4). In addition, we induce a typology of the error cases to systematically organise the vulnerable points of the model (section 6).
8.1 Occlusion based result interpretation
Dataset Misclassified To Change in prediction due to occlusion No change in prediction due to occlusion
FHM Hateful [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@ 8.1 Occlusion based result interpretation
Not-Hateful [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
MAMI Misogynistic [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Not-Misogynistic [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
HARM-C & P Harmful [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Not-Harmful [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Table 4: Occlusion based predictions. The occlusion is implemented by making a given superpixel white. Note: We have hidden explicit nudity in the memes wherever present using black boxes.
Using the SLIC algorithm Süsstrunk et al. (2012) we first segment the misclassfied memes into superpixels. The algorithm automatically segments the images into 5 - 12 superpixels depending on the size of the image. We control the size of each superpixel so that it is neither too small nor too big. Next the region circumscribing each of these superpixels are occluded one at a time by white patches and the model (i.e., LLaVA-1.5 13B with definition and OCR text as input and vanilla as output) is queried again for its predictions. We present a case-by-case manual analysis of the outputs obtained.
@@
CASE 1: Original meme misclassified as positive (i.e., hateful, misogynistic or harmful corresponding to the dataset) and at least one occluded version resulted in the correct prediction (i.e., negative). Table 4 presents some representative examples for this case from each dataset. FHM dataset: Majority of the memes are made up of two images stacked together. These memes put humans and animals (apes/gorilla/goat) in the same frame. Further certain memes have embedded text containing profane words. However, in most cases the overall theme of the meme is not hateful. Occlusion resulted in correct predictions due to the removal of these confusing regions from the meme where the model was misfocusing. That said, our manual inspection indicates that some memes are indeed wrongly annotated as not hateful and the predictions of the model for the original meme are arguably correct. MAMI dataset: Majority of the memes contain perturbed faces of women with weird makeups or portray men either with (i) women or with (ii) embedded text containing words like ‘women’, ‘girlfriend’, ‘girl’. Moreover, many memes are made up of multiple images stacked together. However, the overall theme of the meme is not misogynistic. When occlusion removes the perturbed faces of women or words from the embedded text, the focus of the model is no longer misdirected thus leading to correct predictions. HARM P+C dataset: Here again most of the memes are composed of stacked images. Further many of these memes have long text with small font size embedded on them. Such images are even hard for human judges to label. Owing to this very complex nature of the memes, there in no regular pattern indicating why occluding certain parts of the image results in the correct prediction. This is one case where the occlusion based prediction changes are insufficient in explaining the performance gap of the models and more research is needed in the future.
@@
CASE 2: Original meme misclassified as positive and none of the occluded versions resulted in the correct prediction (see Table 4 for some representative examples). FHM dataset: Surprisingly, we find that a major portion of the memes are indeed hateful and seem to be incorrectly annotated as not hateful. Common targets include religion, gender, race and politicians. Amongst religion, ‘Islam’ is mostly targeted while ‘Hitler’ and ‘Trump’ are the most targeted politicians. None of the occlusions resulted in a change in the predictions which further reinforces the possibility that the data might be wrongly annotated. MAMI dataset: Majority of the memes pose nudity, vulgarity, feminism amongst other attacks on women. Embedded texts have vulgar words like ‘bra’, ‘va**na’, ‘t*ts’, ‘s*xy’, ‘a*s’ targeting women. These memes indeed portray explicit misogyny and as per our analysis, model correctly classifies it as misogynistic and this decision does not get reverted due to occlusion. Here again, we conclude that annotations themselves are incorrect. HARM P+C dataset: Here too we manually observe that most of the memes are indeed harmful and are possibly incorrectly annotated. The predictions of the model seem to be correct and occlusions do not change the predictions.
@@
CASE 3: Original meme misclassified as negative and at least one occluded version resulted in the correct prediction (i.e., positive). Please see Table 4 for some representative examples. FHM dataset: In this group, most of the memes have very small font size of the embedded text. Further the image has multiple objects or numerous color variations. This confuses the model leading to wrong predictions. Occlusion of these confusing regions allowed the model to focus on the parts of the image important for correct classification. MAMI dataset: In most of the cases, image portrays nudity or other forms of vulgarity. In some memes, the embedded text contains the word ‘MILF’ targeting women. Occlusion brings the focus of the model to these disturbing elements of the image leading to the correct prediction. HARM P+C dataset: Majority of the memes contain the image of ‘Trump’ or mention the words ‘Trump’, ‘Covid-19’ or ‘Corona’. Length of embedded texts are very large in these memes which possibly confuses the model. Occlusion helps to bring back the focus of the model to the correct regions resulting in correct predictions.

### v3

@@ Results
vn vn 54.14 38.59 46.86 31.91 44.25 40.16 43.98 33.42 60.94 39.23 51.76 50.03
def vn 51.12 34.55 49.74 34.44 49.12 48.65 48.63 41.05 61.49 42.44 50.12 41.04
ocr vn 50.1 33.73 48.37 33.94 65.44 59.86 48.48 46.96 62.55 38.48 50.4 34.22
def + ocr vn 50.21 34.87 51.49 38.19 64.13 52.84 51.49 44.12 62.77 38.92 51.32 37.32
vn ex 48.38 38.06 50.35 35.2 46.63 41.27 44.84 44.21 38.78 32.18 50.2 34.13
def ex 49.68 33.19 51.43 49.85 46.88 46.77 50.35 50.27 62.46 40.35 49.71 34.67
ocr ex 49.12 34.41 47.39 47.34 65.42 55.19 49.5 45.17 61.66 39.6 49.89 37.13
def + ocr ex 53.06 44.37 54.39 52.52 65.6 51.25 54.09 49.82 61.57 40.84 50.33 35.37
InstructBLIP Flan-T5-xl
vn vn 50.2 33.78 56.67 48.6 64.12 41.88 51.27 36.32 62.68 38.53 50.1 33.73
def vn 50.2 35.13 59.9 56.67 64.97 39.38 51.27 33.89 62.82 38.94 52.1 38
ocr vn 51.2 39.22 55.9 47.75 65.16 40.21 51.27 33.89 62.39 39.12 51.9 37.89
def + ocr vn 52.6 42.33 52.1 39.59 65.54 42.52 51.98 36.68 62.62 39.21 52.2 38.3
vn ex 60.78 60.76 55.34 50.7 60.19 41.53 49.12 46.76 41.39 37.21 48.33 33.14
def ex 51 40.86 53.14 34.92 64.2 39.1 51.46 33.98 36.87 29.08 50.1 33.74
ocr ex 60 58.17 61.13 59.39 61.7 38.16 49.33 33.03 41.47 36.08 50 33.33
def + ocr ex 57.94 55.58 55.01 45.67 65.04 39.41 51.56 34.02 62.61 47.97 49.9 33.29
GPT-4o
vn vn 69.2 68.02 79.3 78.84 71.47 69.54 67.04 66.99 64.84 59.07 65 64.24
def vn 70.4 69.32 79.5 78.95 72.32 71.71 63.66 62.52 62.31 55.56 62.6 61.64
ocr vn 70 69.5 83.6 83.54 67.8 63.46 64.23 63.44 66.95 57.87 61 58.74
def + ocr vn 72.2 71.16 80.3 79.96 71.47 70.86 62.54 62.15 66.1 56.94 61.6 59.57
vn ex 70.4 69.74 82.9 82.78 69.77 66.21 69.3 69.29 65.26 56.01 60.8 59.17
def ex 72.8 72.35 82.03 81.86 72.03 71.06 65.35 64.59 60.76 55.22 64.2 63.5
ocr ex 69.8 69.68 83.03 83.03 69.49 62.82 64.51 62.66 65.54 57.83 61.4 59.47
def + ocr ex 73.4 73.1 82.54 82.52 74.58 72.82 63.94 63.91 65.96 59 61.4 59.47
Table 2: Overall Results: Greyed out cells signify ambiguity above assigned threshold. Best (model, prompt) for each (model, dataset) is underlined. Best (model, prompt) for each dataset is marked bold & underlined; across open source models it is marked bold.
@@
in out ID(9B) LV(13B) LV(7B) IB(V) IV(F) GPT-4o
vn vn 41.11 52.1 38.71 38.07 40.2 68.82
def vn 36.03 57.12 45.85 39.24 42.95 67.69
ocr vn 51.18 53.74 44.16 38.93 41.05 68.36
def + ocr vn 38.78 58.45 52.39 39.86 39.72 68.12
vn ex 38.77 53.35 43.43 36.4 45.44 68.74
def ex 35.26 44.38 42.14 42.94 34.74 69.34
ocr ex 48.81 54.36 41.59 42.94 45.62 68.19
def + ocr ex 37.34 45.84 53.39 45.98 43.93 69.95
Mean 40.91 52.42 45.21 40.54 41.71 68.65
(Std. Dev.) (5.54) (4.65) (4.85) (2.94) (3.37) (0.68)
Table 3: Leaderboard and Stability: Weighted macro-F1 score across 6 datasets with mean + (standard deviation) along promt variants. Overall best macro-F1 score and best mean + (standard deviation) are marked bold and underlined; across open source models are italicized and underlined. Best macro-F1 scores across each prompt variant are underlined; for open source models are italicized. Greyed out cells signify ambiguity above assigned threshold; similar to Table 2. ID(9B): IDEFICS 9B, LV(13B), LV(7B): LLaVA-1.5 13B, 7B, I-BLIP V: InstructBLIP Vicuna 7B, I-BLIP F: InstructBLIP Flan-T5-xl.
@@
In this section, we present the results of our experiments. In Table 2 we show the results for the six datasets across the six models. Each block in the table corresponds to a particular (model, dataset) combination and covers the results for eight prompt pattern combinations. Since we use the generation capability of VLMs for prediction, we observe that in some prompt variants, certain (model, prompt) combinations did not classify the input meme amongst the list_of_labels and diplomatically bypassed the query with an irrelevant answer. This led to a decrease in support to infer the results accurately grounded on correct labels in the dataset. In the table, we have greyed out the cases which did not generate a correct label for at least 90% of the data points. Examples of ambiguous outputs are provided in AppendixD.
@@
Overall results: From Table 2, we observe that InstructBLIP models are not able to correctly predict the labels out of list_of_labels even for English datasets and generate ambiguous answers for quite a large number of prompt variants. Their generated output did not conform with the expected output format specified in input prompt. We also observe that IDEFICS performs best with only OCR as input. LLaVA-1.5 7B, works best with explanation as output, when the input prompt was definition & OCR text. Overall, GPT-4o emerged to be the best model and LLaVA-1.5 13B the best open-source model. While LLaVA-1.5 13B works best with OCR & definition as input and vanilla output, for GPT-4o the best prompt variant is not conclusive as per Table 2. However, we observe that the variation of scores is quite less in GPT-4o compared to open source models and therefore, in the upcoming subsection we formulate a Leaderboard to further understand, quantify and provide in-depth conclusions of our obtained results.
Multilingual capability: We observe that even LLaVA-1.5 7B & 13B models generate ambiguous outputs for BHM and HinGlish datasets; indicating their monolingual limitation. For all models, including GPT-4o, the metrics degrade considerably compared to the results on English datasets. This demonstrates the limitations of VLMs in handling multilingual memes. As expected, GPT-4o outperforms all models in terms of metrics and does not generate any ambiguous output.
@@
Leaderboard: Since engineering solutions are always in the ‘quest for the best’, we propose a quantitative metric to organize the (model, prompt) combinations into a leaderboard. The idea is that the top combinations on this leaderboard should generalize well across the six datasets combined (and eventually across 4 different dimensions of toxicity, i.e., hate, misogyny, harm and offense and 3 different languages, i.e., English, Bangla and HinGlish). For each prompt variant considered over all models, we calculate a weighted average macro-F1 score depending on the number of samples in each of the datasets by the formulation: $\frac{\sum\limits_{\mathcal{D}}(f_{\mathcal{D}})*|\mathcal{D}|}{\sum\limits_{\mathcal{D}}|\mathcal{D}|}$. Here $f_{\mathcal{D}}$ is the macro-F1 for the dataset $\mathcal{D}$. After calculating the weighted average macro-F1 score, we also calculate mean and standard deviation of each model across all prompt variants to obtain the overall performance and stability of the models. The results are shown Table 3.
Based on the above results, we conclude GPT-4o to be the best model with definition & OCR as input and explanation as output. LLaVA-1.5 13B is the best open-source model with definition & OCR text as input and vanilla as output. Moreover, GPT-4o is the best model for all prompt variants taken individually and LLaVA-1.5 13B outperforms other open-source models in 7 of the 8 prompt variants and lags behind LLaVA-1.5 7B variant for the setting with definition & OCR as input and explanation as output.
Stability: Further, we also observe that GPT-4o is not only the best model across all prompt variants, but is also the most stable model with least standard deviation across different prompt variants. Amongst open-source models, LLaVA-1.5 13B model has the best mean.
Baselines: As per the discussion in ‘Models’ section, we present the results in Table 4 on zero-shot and FHM fine-tuned versions of the baseline. Note that it is done to be very near to zero-shot evaluation. GPT-4o outperforms these baselines even without any fine-tuning. Notably, LLaVA-1.5 13B version also outperforms these baselines across Macro-F1 score. This raises a deep concern about the generalization capability of previous works.
Datasets Metrics zero shot fine-tuned
FHM acc 52.13 (0.95) 60.67 (1.33)
mf1 47.11 (0.97) 57.27 (1.53)
MAMI acc 50.43 (0.51) 59.57 (0.32)
mf1 43.88 (0.47) 56.58 (0.43)
HARM-C acc 57.72 (0.99) 61.21 (1.27)
mf1 46.57 (1.52) 44.72 (1.34)
HARM-P acc 52.14 (1.73) 54.76 (0.65)
mf1 44.67 (2.65) 44.81 (1.51)
BHM acc 57.48 (1.38) 60.62 (1.02)
mf1 44.92 (2.1) 50.28 (1.73)
HinGlish acc 50.2 (0.6) 55.13 (0.23)
mf1 38.55 (0.68) 49.03 (0.41)
Table 4: Baselines: Mean with standard deviation in (parenthesis). We have discussed details in Section ‘Models’.
Error analysis
@@
In the previous section we concluded that GPT-4o (with definition + OCR text as input and explanation as output) is the best model and LLaVA-1.5 13B (with definition + OCR text as input and vanilla as output) is the best model amongst chosen open source models. Since it is practically infeasible to study error analysis over all models and prompt variants, we therefore investigate the cases of misclassification for these two best models (one proprietary and one open source based on the Leaderboard) with their best prompt variants. We comprehensively evaluate GPT-4o for a total of 828 misclassified memes: 124- FHM, 149- MAMI, 195- HARM P+C, 205- BHM and 155- HinGlish datasets and LLaVA-1.5 13B for a total of 1184 misclassified memes; 202- FHM, 321- MAMI, 276- HARM P+C, 216- BHM and 169- HinGlish datasets. In particular we attempt to obtain an explanation of parts in the meme that confuses the model and leads to mispredictions (first sub-section). Further, we induce a typology of the error cases to systematically organise the vulnerable points of the model (second sub-section).
Occlusion based result interpretation
Dataset Model CASES: 1, 2 CASES: 3, 4
FHM GPT-4o 15.32, 55.65 8.06, 20.97
LV-13B 14.36, 26.73 20.3, 38.61
MAMI GPT-4o 17.45, 42.95 14.77, 24.83
LV-13B 20.87, 15.89 38.32, 24.92
HARM P+C GPT-4o 18.97, 33.85 12.82, 34.36
LV-13B 34.78, 30.07 20.29, 14.86
BHM GPT-4o 12.68, 9.76 26.83, 50.73
LV-13B 7.87, 2.31 35.19, 54.63
HinGlish GPT-4o 13.55, 7.1 34.19, 45.16
LV-13B 18.93, 5.33 42.60, 33.14
Table 5: Case Distribution and Rigidness: Percentage distribution of misclassified samples from each dataset for all 4 CASES studied under occlusion based interpretability. All numbers are in percentage (%). Wherever CASE 2 or 4 are higher than CASE 1 or 3 respectively, the entries are marked in bold as a measure of rigidness of models towards occlusion.
Dataset Model CASE 1 CASE 2 CASE 3 CASE 4
FHM GPT-4o
NHM: 47.37
WA: 36.84
SI: 26.32
WA:72.46
NHM: 18.84
MCVG: 50
URET: 30
IM: 73.08
LV-13B
WA: 41.38
SI: 24.14
NHM: 24.14
WA: 62.96
NHM: 24.07
MCVG: 41.46
URET: 19.51
SI: 12.2
IM: 74.36
MAMI GPT-4o
WA: 69.23
PFW: 34.62
WA: 64.06 NV: 54.54 IM: 70.27
LV-13B
WA: 44.78
SI: 19.4
PFW: 17.91
WA: 64.71 NV: 47.15 IM: 56.25
HARM CP GPT-4o No conclusion WA: 68.18 URET: 24 FC: 26.67
LV-13B No conclusion WA: 40.34 URET: 30.36 FC: 29.27
BHM GPT-4o
SI: 42.56
SI: 55 SI: 60 SI: 50.77
LV-13B
SI: 44.89
SI: 69.65 SI: 65.89 SI: 59.7
HinGlish GPT-4o
SI: 57.14
SI: 63.64 No conclusion WA: 55.5
LV-13B
SI: 54.54
SI: 62.51 No conclusion WA: 56.25
Table 6: Class Based Error Typology: Distribution of various classes across misclassifications for GPT-4o & LLaVA-1.5 13B as per manual evaluation done on occlusion. All the numbers are in percentage (%). Following are the 9 classes covered: NHM: Not hateful memes containing (i) common target words (i.e., Islam, white, migrants, etc.) (ii) image signifying these common targets or having some politician in it (iii) profane words, WA: Wrong annotation, SI: Stacked images, MCVG: Multiple color variations, grayscale image or multiple objects, URET: Lengthy, unreadable embedded text or having small font size, IM: Implicitly hate meme, PFW: Perturbed / animated faces of women, NV: Nudity or vulgarity in image or embedded text, FC: Fake conversation. Note: HARM CP: Harm C+P, LV-13B: LLaVA-1.5 13B.
Using the SLIC algorithm (Süsstrunk et al. 2012) [^9] Useful tutorial on SLIC algorithm -
https://darshita1405.medium.com/superpixels-and-slic-6b2d8a6e4f08
we first segment the misclassfied memes into superpixels. The algorithm automatically segments the images into 5 - 12 superpixels depending on the size of the image. We control the size of each superpixel so that it is neither too small nor too big. Next the region circumscribing each of these superpixels are occluded one at a time by white patches and the models (i.e., GPT-4o and LLaVA-1.5 13B) are queried again for its predictions with their corresponding best prompt variant (as per Leaderboard).
CASES: Following are the cases in which we have broadly divided our occlusion study: CASE 1- Original meme misclassified as positive (i.e., hateful, misogynistic, harmful, or offensive corresponding to the dataset) and at least one occluded version resulted in the correct prediction (i.e., negative). CASE 2- Original meme misclassified as positive and none of the occluded versions resulted in the correct prediction . CASE 3- Original meme misclassified as negative and at least one occluded version resulted in the correct prediction (i.e., positive). CASE 4- Original meme misclassified as negative and none of the occluded versions resulted in the correct prediction.
Rigidness: In Table 5, we present the distribution for each of the cases. For GPT-4o, we observe that CASE 2 & CASE 4 have higher values compared to CASE 1 and CASE 3 respectively. This points to a very interesting observation regarding the rigidness of GPT-4o in its classification decision. LLaVA-1.5 13B on the other hand has similar behaviour for the FHM and BHM datasets. Overall these results show that the models which perform better are also more rigid in their generation and are comparatively less prone to a decision change based on the perturbations in the image. Decrease in rigidness on non-English datasets relates with the weaker performance of models on multilingual ability.
Case by case study: We perform further analysis and divide each case across 9 classes (refer Table 6) and provide representative examples in Table 7. Based on these tables we present a case-by-case manual analysis of the results. We employ three annotators to perform manual analysis of these occluded images. Out of them, two are NLP researchers having professional working experience on understanding hateful memes and the other one is an undergraduate student in his fourth year with major in Computer Science and has working experience in NLP. All the three annotators are from the authors’ institute and are well aware of the topicality of discussion being conveyed in the memes. As we employ three annotators, we calculate Krippendorff’s $\alpha$ (Castro 2017) as a metric for inter-annotator agreement. We obtain high agreement values of 0.873 & 0.875 for GPT-4o & LLaVA-1.5 13B, respectively. Note that we have added annotation guidlines in Appendix.
Dataset Misclassified to Change in prediction No change in prediction
FHM Hateful [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Not-Hateful [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
MAMI Misogynistic [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Not-Misogynistic [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
HARM-C & P Harmful [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Not-Harmful [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Table 7: Case Studies: Occlusion based predictions. The occlusion is implemented by making a given superpixel white. For each case, we present one sample each from GPT-4o & LLaVA-1.5 13B, respectively. For CASE 1 and CASE 3 we provide both the occluded and non-occluded samples. Due to space constraints and no robust conclusion from manual annotation of BHM and HinGlish, we omit them.
@@
CASE 1: FHM dataset: Around one-fourth of the memes are made up of multiple images stacked together. These memes put humans and animals (apes/gorilla/goat) in the same frame. Further, certain not hateful memes also seem to contain profane words or mention the common target words either as embedded text or as a representation in the image itself (refer Table 6’s caption for further details). Occlusion results in correct predictions due to the removal of these confusing regions from the meme where the model was misfocusing. That said, our manual inspection indicates that there are memes which are indeed wrongly annotated as not hateful and the predictions of the model for the original meme are arguably correct. MAMI dataset: Majority of the memes are indeed wrongly annotated or contain perturbed faces of women with weird makeups or portray men either with (i) women or with (ii) embedded text containing words like ‘women’, ‘girlfriend’, ‘girl’. Moreover, for LLaVA-1.5 13B, some memes are made up of multiple images stacked together. However, the overall theme of the meme is not misogynistic. When occlusion removes the perturbed faces of women or words from the embedded text, the focus of the model is no longer misdirected thus leading to correct predictions. HARM P+C dataset: Here again most of the memes are composed of stacked images. Further many of these memes have long text with small font size embedded on them. Such images are even hard for human judges to label. Owing to this very complex nature of the memes, there in no regular pattern indicating why occluding certain parts of the image results in the correct prediction. This is a case where the occlusion based prediction changes are insufficient in explaining the performance gap of the models and more research is needed in the future. BHM + HinGlish datasets: No conclusive pattern was visible due to the diverse topicality of discussion and outdated conversations within current context. However, majority of the memes were composed of stacked images.
@@
CASE 2: FHM dataset: Surprisingly, we find that a major portion of the memes are indeed hateful and seem to be incorrectly annotated as not hateful. Common targets include religion, gender, race and politicians. Amongst religion, ‘Islam’ is mostly targeted while ‘Hitler’ and ‘Trump’ are the most targeted politicians. None of the occlusions resulted in a change in the predictions which further reinforces the possibility that the data might be wrongly annotated. MAMI dataset: Majority of the memes pose nudity, vulgarity, feminism amongst other attacks on women. Embedded texts have vulgar words like ‘bra’, ‘va**na’, ‘t*ts’, ‘s*xy’, ‘a*s’ targeting women. These memes indeed portray explicit misogyny and as per our analysis, model correctly classifies it as misogynistic and this decision does not get reverted due to occlusion. Here again, we conclude that annotations themselves are incorrect. HARM P+C dataset: Here too we manually observe that most of the memes are indeed harmful and are possibly incorrectly annotated. The predictions of the model seem to be correct and occlusions do not change the predictions. BHM + HinGlish datasets: Similar to CASE 1, no conclusive pattern was visible and majority of the memes were composed of stacked images.
@@
CASE 3: FHM dataset: In this group, most of the memes have very small font size of the embedded text. Further the image has multiple objects or numerous color variations. This confuses the model leading to wrong predictions. Occlusion of these confusing regions allowed the model to focus on the parts of the image important for correct classification. MAMI dataset: In most of the cases, image portrays nudity or other forms of vulgarity. In some memes, the embedded text contains the words like ‘MILF’, ‘VIRGIN’, etc.. targeting women. Occlusion brings the focus of the model to these disturbing elements of the image leading to the correct prediction. HARM P+C dataset: Majority of the memes contain the image of ‘Trump’ or mention the words ‘Trump’, ‘Covid-19’ or ‘Corona’. Length of embedded texts are very large, which possibly confuses the model. Occlusion helps to bring back the focus of the model to the correct regions resulting in correct predictions. BHM + HinGlish datasets: While no conclusion was visible for HinGlish, in BHM majority of the memes were composed of stacked images.
Figure 2: Typology: Green circles represent misclassification to positive label; red signifies misclassification to negative label. Each set of misclassification is bifurcated into two clusters. Distribution of cases, topic words and representative image cluster are shown for GPT-4o and LLaVA-1.5 13B models, for FHM, MAMI and HARM-C + P datasets. Enlarged image clusters and results for BHM and HinGlish are in Appendix. Important keywords in each topic are marked in bold.